OpenStudy (anonymous):

solve systems of equations. determine if there are two circles or a circle and a line. {x^2+y^2=4 [x^2+2x+y^2=0

OpenStudy (anonymous):

alright since we know x^2+y^2=4 let us sub this into our second equation x^2+2x+y^2=0 to get 2x +4 =0 and x=-2 and now let us sub this value into any of the equation to get y =0 and that is our intersection points and these are two circles let me solve the second equation we have x^2+2x+y^2=0 and that is just x^2+2x +1 - 1+y^2=0 --> (x+1)^2 +y^2 = 1 and now let me draw this for you :)

OpenStudy (anonymous):



OpenStudy (anonymous):

did you get what i did @tylerjh

OpenStudy (anonymous):

uhmmmkinda

OpenStudy (anonymous):

ok iunderstand so its two circles and wats the final answer?